OpenStudy (loser66):

Converge or diverge? \[\sum_{n=1}^\infty \dfrac{log(n+1)-log n}{tan^{-1}(2/n)}\] Please, help

OpenStudy (amistre64):

im wondering if a ratio test give us anything conclusive? it appears the top goes to 0, and the bottom goes to 0 might be lhopable?

OpenStudy (loser66):

Apply l'hospital: the summand is numerator: (1/(n+1) - (1/n) = -1/(n^2+n) denominator : (1/1+4n^2) combine: \[-\dfrac{1+4n^2}{n^2+n}\] hence the lim =-4 \(\neq 0\) divergence, right?

OpenStudy (amistre64):

ill have to look that up to be sure; but memory wants to say: if the limit doesnt go to zero for the terms, it diverges. http://www.mathwords.com/l/limit_test_for_divergence.htm

OpenStudy (amistre64):

and by the limit test, it diverges .. as we expected.

OpenStudy (loser66):

Oh yea, I messed up. thank you.